OpenStudy (anonymous):

I need some help with this question on conics: Find an equation for the ellipse with the foci (1,1) and (-1,-1) and a major axes of length 4.

OpenStudy (anonymous):

This is a shifted conic, i need to find the equation of this shifted conic

OpenStudy (dumbcow):

this looks like an ellipse with a slanted major axis, is that correct?

OpenStudy (anonymous):

yes thats correct

OpenStudy (anonymous):

It lies on the rotated XY coordinate axes

OpenStudy (dumbcow):

ok the axis is the line y=x, since it goes grough both foci points 2a = 4 a = 2 imagine a triangle with equal legs and hypotenuse of 2 x =y = 2/sqrt2 = sqrt2 verticis are (sqrt2, sqrt2), (-sqrt2, -sqrt2)

OpenStudy (dumbcow):

minor axis must be perpendicular so on the line y=-x length of minor axis is 2b, c is length of foci point from origin --> c=sqrt2 a^2 = b^2 +c^2 2^2 = b^2 +(sqrt2)^2 4 = b^2 +2 b = sqrt2 again imagine 45-45 -90 triangle with hypotenuse of sqrt2 x=y=1 minor verticis = (-1,1) (1,-1)

OpenStudy (anonymous):

Shoudlnt we incorporate those formulas x=Xcos(theta)-Ysin(theta) and y=Xsin(theta)+Ycos(theta)

OpenStudy (anonymous):

because i am looking for the equation of the rotated ellipse

OpenStudy (dumbcow):

yes to get the equation we have to rotate 45 degrees to get major axis on line y=x now we have a,b so equation of non-rotated ellipse is: \[\frac{x^{2}}{4}+\frac{y^{2}}{2} = 1\] theta = 45 sin(45) = cos(45) = sqrt2/2 \[x' = \sqrt{2}/2(x-y)\] \[y' = \sqrt{2}/2 (x+y)\] replace x with x' and y with y' in equation

OpenStudy (anonymous):

Okay i got the equation of the non rotated ellipse, but now i understand how to get the equation of the rotated ellipse. Thank you.

OpenStudy (anonymous):

if you could take a look at my other question regarding a helix and curvature, it would be much appreaciated

OpenStudy (dumbcow):

ok